A rhombus has diagonals of 32 and 20 inches. Find the angle of opposite the longer diagonal
1
Expert's answer
2019-07-25T14:24:13-0400

Rhombus side

"a=\sqrt{16^2+10^2}=\sqrt{356},"

"angle=2\arcsin{\frac{16}{a}}=2\arcsin{\frac{16}{\sqrt{356}}}\approx 116^{\circ}."

Answer: the angle is approximately equal to 116 degrees.